### v2

# Hölder spiral arcs

Abstract.
We establish a quantitative necessary and sufficient condition for a spiral arc to be a Hölder arc. The class of spiral arcs contains spirals studied by Fraser in [Fra21], and by Burell-Falconer-Fraser in [BFF22]. As an application, we recover the sharp result on the Hölder winding problem for polynomial spirals, initially proved in [Fra21]. Moreover, we provide a sharp exponent estimate for the Hölder classification of polynomial spirals, which coincides with the corresponding quasiconformal classification estimate, and improve certain exponent bounds on the Hölder classification of elliptical spirals from [BFF22].
Key words and phrases:
Spiral, Hölder arc, Hölder exponents
2020 Mathematics Subject Classification:
Primary 26A16; Secondary 28A80
V.V was partially supported by NSF DMS grant 2154918.
1. Introduction
Given a function $\phi:[2\pi,+\infty)\to(0,\infty)$ with $\lim_{t\rightarrow\infty}\phi(t)=0$, we denote by $\mathcal{S}_{\phi}$ the spiral
$\{\phi(t)e^{\mathbf{i}t}:t\in[2\pi,+\infty)\}\cup\{(0,0)\}.$
Spirals hold a prominent role in fluid turbulence [FHT01, VH91, FHV93], dynamical systems [ZZ05, HVZZ23], and even certain types of models in mathematical biology [TAMM89, Mur02]. Moreover, they provide examples of “non-intuitive” fractal behavior (see [DMFT83]), while they have also been extensively studied due to their unexpected analytic properties. For instance, Katznelson-Nag-Sullivan [KNS90] demonstrated the dual nature of spirals $\mathcal{S}_{\phi}$ for decreasing $\phi$, lying in-between smoothness and “roughness”, as well as their connection to certain Riemann mapping questions. The existence of Lipschitz and Hölder parametrizations of certain $\mathcal{S}_{\phi}$ has also been studied by the aforementioned authors in [KNS90], by Fish-Paunescu in [FP18], and by Fraser in [Fra21].
In particular, Fraser in [Fra21] focuses on polynomial spirals where $\phi(t)=t^{-p}$, for $p>0$, and shows that $\mathcal{S}_{p}:=\mathcal{S}_{\phi}$ is an $\alpha$-Hölder arc for all $\alpha\in(0,p)$, with this upper bound on the exponent $\alpha$ being sharp. In the same paper, Fraser suggests a programme of research focused on determining quantitative conditions under which two sets are bi-Hölder equivalent (see [Fra21, p. 3254]). Towards this direction, Burell-Falconer-Fraser [BFF22] further studied the elliptical spirals
$\mathcal{S}_{p,q}=\{t^{-p}\cos t+\mathbf{i}t^{-q}\sin t:t\in[2\pi,\infty)\}\cup\{(0,0)\},$
and provided bounds on the exponent of Hölder maps between two such spirals. Note that all elliptical spirals $\mathcal{S}_{p,q}$ can be written in the form $\mathcal{S}_{\phi}$, for some appropriate function $\phi:[2\pi,+\infty)\to(0,\infty)$.
Motivated by the interest in the regularity and Hölder classification of continuous spirals with no self-intersections, we define and study a general class of spirals that contains those studied in [Fra21, BFF22]. Given a continuous $\phi:[2\pi,+\infty)\to(0,\infty)$ with $\lim_{t\rightarrow\infty}\phi(t)=0$, for all $j\in\mathbb{N}$ set
$\mathcal{S}_{\phi}^{j}:=\{\phi(t)e^{\mathbf{i}t}:t\in[2\pi j,2\pi(j+1)]\},$
and
$\phi_{j}:=\max\{\phi(t):t\in[2\pi j,2\pi(j+1)]\}.$
We say $\mathcal{S}_{\phi}$ is almost circular if there is $C_{\phi}>0$ such that $\ell(\mathcal{S}_{\phi}^{j})\leq C_{\phi}\phi_{j}$, for all $j\in\mathbb{N}$, where $\ell(\mathcal{S}_{\phi}^{j})$ denotes the length of $\mathcal{S}_{\phi}^{j}$. The main result is a necessary and sufficient quantitative condition that a spiral arc (i.e. spiral with no self-intersections) with the above property needs to satisfy in order to be a Hölder arc.
Theorem 1.1.
Let $s>1$ and $\mathcal{S}_{\phi}$ be an almost circular spiral arc. Then $\mathcal{S}_{\phi}$ is a $(1/s)$-Hölder arc if, and only if, $\sum_{n=1}^{\infty}\phi_{n}^{s}$ converges.
In particular, Theorem 1.1 follows from an even more general result we prove for all spiral arcs $\mathcal{S}_{\phi}$ with partitions $\mathcal{S}_{\phi}^{j}$ that are Hölder in a uniform way; see Theorem 3.1. The almost circular property is introduced for mainly two reasons. First, quantitative conditions on the Hölder regularity of spirals of the form $\mathcal{S}_{\phi}$ would be at least as difficult to establish as those for graphs of functions $\phi$, which is generally a challenging problem. Second, and as already mentioned, the polynomial spiral arcs from [Fra21] and the elliptical spiral arcs from [BFF22] are in fact almost circular, thus generalizing these already interesting classes. As a result, Theorem 1.1 allows us to recover the sharp exponent for the Hölder winding problem studied in [Fra21], and to improve the results on the Hölder classification of spirals $\mathcal{S}_{p,q}$ that were previously established in [BFF22] (see Section 4 for details). In fact, in the case of spirals of the form $\mathcal{S}_{p}$, $p>0$, Theorem 1.1 provides a sharp estimate on their Hölder classification in the following sense.
Theorem 1.2.
Let $0<r\leq p$. There is a $r/p$-Hölder map $f:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$, and a Lipschitz map $g:\mathcal{S}_{r}\rightarrow\mathcal{S}_{p}$. Moreover, every $\alpha$-Hölder map $h:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$ needs to satisfy $\alpha\leq r/p$.
This paper is organized as follows. In Section 2 a characterization of a Hölder arc is established by using the notion of variation of a metric arc. Section 3 contains the proof of Theorem 1.1, which uses the aforementioned characterization. In Section 4 we use Theorem 1.1 to recover the sharp exponents of Hölder regularity for polynomial spirals $\mathcal{S}_{p}$ from [Fra21] and improve the classification estimates for elliptical spirals $\mathcal{S}_{p,q}$ from [BFF22]. In the same section, the proof of Theorem 1.2 and further remarks on the relation between the Hölder and the quasiconformal classification problem of spirals are included.
Background and notation
Let $(X,d_{X}),(Y,d_{Y})$ be non-empty metric spaces. We say that a map $f:X\rightarrow Y$ is $\alpha$-Hölder (continuous), for some $\alpha\in(0,1)$, if there is $C>0$ such that
$d_{Y}(f(x_{1}),f(x_{2}))\leq Cd_{X}(x_{1},x_{2})^{\alpha},$
for all $x_{1},x_{2}\in X$. The number $\alpha$ is called the Hölder exponent of $f$, and the smallest $C>0$ is the Hölder semi-norm of $f$, denoted by $\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{\alpha}f$.
Recall that a metric space $X$ is a metric arc if there is a homeomorphism $f$ mapping the interval $[0,1]$ onto the space $X$. Given an interval $I\subset[0,1]$, we say that $f(I)$ is a subarc of $X$. Furthermore, if the interval $I$ has endpoints $a,b\in[0,1]$, we say that $f(I\setminus\{a,b\})$ is the interior of the subarc $f(I)$. If $X$ and $f(I)$ have distinct interiors, we say that $f(I)$ is a proper subarc of $X$. If $f$ is $\alpha$-Hölder for some $\alpha\in(0,1)$, we say that $X$ is an $\alpha$-Hölder arc.
For $s>0$, $\epsilon>0$, and a subset $E$ of the metric space $X$, the $s$-dimensional $\epsilon$-approximate Hausdorff measure of $E$ is defined as
$\mathcal{H}_{\epsilon}^{s}(E)=\inf\left\{\sum_{i}(\operatorname{diam}U_{i})^{s}:\{U_{i}\}\text{\ countable cover of}\,\,E\,\,\text{with}\,\,\operatorname{diam}U_{i}\leq\epsilon\right\}.$
The $s$-dimensional Hausdorff (outer) measure of $E$ is the limit
$\mathcal{H}^{s}(E)=\lim_{\epsilon\to 0}\mathcal{H}_{\epsilon}^{s}(E).$
2. The $s$-variation and $(1/s)$-Hölder rectifiability
A classification of Lipschitz arcs (and even more general Lipschitz curves) was given by Ważewski in 1927: an arc is a Lipschitz arc if and only if $\mathcal{H}^{1}(X)<\infty$, and if $X$ is a Lipschitz arc, then there exists a $(2\mathcal{H}^{1}(X))$-Lipschitz parameterization $f:[0,1]\to X$. See [AO17, Theorem 4.4] for the proof. An analogous result for $\frac{1}{s}$-Hölder arcs does not exist as there is no connection between $\frac{1}{s}$-Hölder parameterizations and $\mathcal{H}^{s}$; see for example [BNV19, Proposition 9.6]. In this section we give a characterization of $\frac{1}{s}$-Hölder arcs in terms of the $s$-variation, the correct analogue of $\mathcal{H}^{1}$ for $s\geq 1$.
A partition of $X$ is a finite collection of subarcs $\mathcal{P}=\{X_{1},\dots,X_{n}\}$ with disjoint interiors, and with their union being equal to $X$. For $s\geq 1$, define the $s$-variation of $X$ by
(2.1) $\|X\|_{s\text{-var}}:=\sup_{\mathcal{P}}\sum_{X^{\prime}\in\mathcal{P}}(\operatorname{diam}{X^{\prime}})^{s}\in[0,+\infty],$
with the supremum being over all partitions of $X$.
For $s\geq 1$, we set $H_{1/s}(X)$ to be the infimum of all constants $H>0$ for which there is a surjection $f:[0,1]\to X$ satisfying
$|f(x)-f(y)|\leq H|x-y|^{1/s},\quad\text{for all $x,y\in[0,1]$.}$
If no such $f$ exists, then $H_{1/s}(X)=\infty$.
The relation between quantities $\|X\|_{s\text{-var}}$ and $H_{1/s}(X)$, and the existence of Hölder parameterizations is given in the next proposition.
Proposition 2.1.
Let $X$ be a metric arc and $s\geq 1$. Then
$\|X\|_{s\text{-var}}=H_{1/s}(X)^{s}.$
In particular, $X$ is a $(1/s)$-Hölder arc if, and only if, $\|X\|_{s\textup{-var}}<\infty$.
The Hölder regularity of a metric arc $X$ has been closely tied before to variation notions defined for continuous maps $g:[0,1]\rightarrow X$ (see [FV10, Definitions 1.1, 5.1]). Specifically, given a homeomorphism $f:[0,1]\rightarrow X$, the methods leading to the proof of [FV10, Proposition 5.14] could be applied to $f$ and yield Proposition 2.1. However, it should be noted that the statement of [FV10, Proposition 5.14] is dependent on each given continuous $g:[0,1]\rightarrow X$. Since our definition of $s$-variation is intrinsically more geometric and does not depend on any given parametrization of $X$, we include the proof in this context.
For the proof of Proposition 2.1 we require several lemmas.
Lemma 2.2.
Let $X$ be a metric arc and $s>1$.
  1. (1)
    If $X^{\prime}$ is a subarc of $X$, then $\|X^{\prime}\|_{s\textup{-var}}\leq\|X\|_{s\textup{-var}}$.
  2. (2)
    If $\|X\|_{s\textup{-var}}<\infty$ and $X^{\prime}$ is a proper subarc of $X$, then $\|X^{\prime}\|_{s\textup{-var}}<\|X\|_{s\textup{-var}}$.
  3. (3)
    We have $\|X\|_{s\textup{-var}}\geq\max\{\mathcal{H}^{s}(X),(\operatorname{diam}{X})^{s}\}$.
  4. (4)
    If $X=X_{1}\cup\cdots\cup X_{n}$ is a partition of $X$ into subarcs, then
    $\|X\|_{s\textup{-var}}\geq\|X_{1}\|_{s\textup{-var}}+\cdots+\|X_{n}\|_{s\textup{-var}}.$
Proof.
Property (1) is immediate from the definition.
For (2), assume that $\|X\|_{s\textup{-var}}<\infty$ and that $X^{\prime}$ is a proper subarc of $X$. Let $Y$ be a subarc of $X$ that intersects with $X^{\prime}$ only at an endpoint. Let also $X_{1},\dots,X_{n}$ be a partition of $X^{\prime}$ such that
$\|X^{\prime}\|_{s\textup{-var}}\leq\sum_{i=1}^{n}(\operatorname{diam}{X_{i}})^{s}+\tfrac{1}{2}(\operatorname{diam}{Y})^{s}.$
Then,
$\|X\|_{s\textup{-var}}\geq\|X^{\prime}\cup Y\|_{s\textup{-var}}\geq\sum_{i=1}^{n}(\operatorname{diam}{X_{i}})^{s}+(\operatorname{diam}{Y})^{s}>\|X^{\prime}\|_{s\textup{-var}}.$
For (3), note first that $\mathcal{P}=\{X\}$ is a partition of $X$ so
$\|X\|_{s\text{-var}}\geq(\operatorname{diam}{X})^{s}.$
To show that $\|X\|_{s\text{-var}}\geq\mathcal{H}^{s}(X)$, fix $\epsilon>0$ and let $\mathcal{P}$ be a partition of $X$ such that $\operatorname{diam}{X^{\prime}}<\epsilon$ for all $X^{\prime}\in\mathcal{P}$. Then,
$\mathcal{H}^{s}_{\epsilon}(X)\leq\sum_{X^{\prime}\in\mathcal{P}}(\operatorname{diam}{X^{\prime}})^{s}\leq\|X\|_{s\text{-var}}.$
Letting $\epsilon$ go to 00, we obtain the desired inequality.
The proof of (4) follows immediately from the definition of $s$-variation. ∎
Lemma 2.3.
Let $s\geq 1$, let $X$ be a metric arc with $\|X\|_{s\textup{-var}}<\infty$, and let $f:[0,1]\to X$ be a homeomorphism. Then the function $t\mapsto\|f([0,t])\|_{s\textup{-var}}$ is continuous.
Proof.
The continuity of the function $t\mapsto\|f([0,t])\|_{s\textup{-var}}$ follows by [FV10, Proposition 5.8] applied to the given homeomorphism $f$.
We are now ready to prove Proposition 2.1.
Proof of Proposition 2.1.
Suppose first that $H_{1/s}(X)<\infty$. Then, there exists a $(1/s)$-Hölder homeomorphism $f:[0,1]\to X$ with $H_{1/s}(X)\leq\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$. Fix arbitrary $H>\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$ and $\epsilon\in(0,1)$.
Let $X_{1},\dots,X_{n}$ be a partition of $X$ and for each $i\in\{1,\dots,n\}$ let $I_{i}=f^{-1}(X_{i})$. Then $\{I_{1},\dots,I_{n}\}$ is an interval partition of $[0,1]$. For each $i\in\{1,\dots,n\}$, let $x_{i},y_{i}\in I_{i}$ such that
$|f(x_{i})-f(y_{i})|\geq(1-\epsilon)\operatorname{diam}{f(I_{i})}$
and denote by $J_{i}$ the interval in $I_{i}$ with endpoints $x_{i},y_{i}$. By choice of $x_{i},y_{i}$ and Hölder continuity of $f$, we have
$(\operatorname{diam}{f(I_{i})})^{s}\leq(1-\epsilon)^{-s}|f(x_{i})-f(y_{i})|^{s}\leq(1-\epsilon)^{-s}H^{s}\operatorname{diam}{J_{i}}.$
Since $I_{i}=f^{-1}(X_{i})$, the above implies
$\sum_{i=1}^{n}(\operatorname{diam}{X_{i}})^{s}\leq(1-\epsilon)^{-s}H^{s}\sum_{i=1}^{n}\operatorname{diam}{J_{i}}\leq(1-\epsilon)^{-s}H^{s}.$
Taking supremum over all partitions and letting $\epsilon\to 0$ we obtain $\|X\|_{s\text{-var}}\leq H^{s}$. Since $H>\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$ is arbitrary, and $\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{f}$ can be chosen as close to $H_{1/s}(X)$ as necessary, this implies that $\|X\|_{s\text{-var}}\leq H_{1/s}(X)^{s}$.
For the converse, let $f:[0,1]\to X$ be a homeomorphism, $s\geq 1$, and assume that $\|X\|_{s\text{-var}}<\infty$. For each $x\in[0,1]$ define
$\psi(x)=\frac{\|f([0,x])\|_{s\text{-var}}}{\|X\|_{s\text{-var}}}.$
By Lemma 2.2, $\psi$ is an increasing function from $[0,1]$ into $[0,1]$ and, by Lemma 2.3, $\psi$ is continuous. Since $\psi(0)=0$ and $\psi(1)=1$, it follows that $\psi$ is surjective. By Lemma 2.2(2), it follows that $\psi$ is in fact a homeomorphism. This allows for the definition of $F=f\circ\psi^{-1}:[0,1]\to X$.
It remains to show that $F$ is $(1/s)$-Hölder. Let $0\leq x<y\leq 1$ and let $0\leq x^{\prime}<y^{\prime}\leq 1$ be such that $\psi(x^{\prime})=x$ and $\psi(y^{\prime})=y$. Then,
$|F(x)-F(y)|^{s}$ $=|f(x^{\prime})-f(y^{\prime})|^{s}$
$\leq\|f([x^{\prime},y^{\prime}])\|_{s\text{-var}}$
$\leq\|f([0,y^{\prime}])\|_{s\text{-var}}-\|f([0,x^{\prime}])\|_{s\text{-var}}$
$=\|X\|_{s\text{-var}}|\psi(x^{\prime})-\psi(y^{\prime})|$
$=\|X\|_{s\text{-var}}|x-y|.$
Therefore, $F$ is $(1/s)$-Hölder continuous with $\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{F}\leq\|X\|_{s\text{-var}}^{1/s}$. As a result,
$H_{1/s}(X)\leq\mathop{\mathrm{missing}}{H\ddot{o}ld}\nolimits_{1/s}{F}\leq\|X\|_{s\text{-var}}^{1/s}<\infty,$
which allows to apply the first direction and yields $H_{1/s}(X)=\|X\|_{s\text{-var}}^{1/s}$ as needed. This equality and the definition of $H_{1/s}(X)$ complete the proof. ∎
3. Hölder rectifiability of spiral arcs
Fix for the rest of this section a spiral arc $\mathcal{S}_{\phi}$, and recall that for $j\in\mathbb{N}$, we set
$\mathcal{S}_{\phi}^{j}=\{\phi(t)e^{\mathbf{i}t}:t\in[2\pi j,2\pi(j+1)]\}.$
Note that since $\mathcal{S}_{\phi}$ is an arc, the sequence $(\operatorname{diam}\mathcal{S}_{\phi}^{j})_{j\in\mathbb{N}}$ is decreasing. We prove the following result for general spiral arcs.
Theorem 3.1.
Let $s\geq 1$. Then $\mathcal{S}_{\phi}$ is a $\frac{1}{s}$-Hölder arc if and only if $\sum_{j=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}$ converges.
Before proving Theorem 3.1, let us first give the proof of Theorem 1.1.
Proof of Theorem 1.1.
Suppose that $\mathcal{S}_{\phi}$ is almost circular. Fix $j\in\mathbb{N}$ and $s\geq 1$. On the one hand, if $h_{j}:[0,1]\to\mathcal{S}_{\phi}^{j}$ is the constant speed Lipschitz map, then for all $x,y\in[0,1]$,
$|h_{j}(x)-h_{j}(y)|\leq\ell(\mathcal{S}_{\phi}^{j})|x-y|\leq\ell(\mathcal{S}_{\phi}^{j})|x-y|^{1/s}\leq C_{\phi}\phi_{j}|x-y|^{1/s}.$
On the other hand, if $h:[0,1]\to\mathcal{S}_{\phi}^{j}$ is $\frac{1}{s}$-Hölder with Hölder constant $H$, then there exist $x,y\in[0,1]$ such that $|h(x)-h(y)|=\operatorname{diam}{\mathcal{S}_{\phi}^{j}}$ which gives
$H\geq H|x-y|^{1/s}\geq|h(x)-h(y)|=\operatorname{diam}{\mathcal{S}_{\phi}^{j}}\geq\phi_{j}.$
Hence, $H_{1/s}(\mathcal{S}_{\phi}^{j})\simeq\phi_{j}$, and the statement follows directly from Theorem 3.1. ∎
We now turn to the proof of Theorem 3.1.
Proof of Theorem 3.1.
Suppose first that $\mathcal{S}_{\phi}$ is a $(1/s)$-Hölder arc. Fix $k\in\mathbb{N}$, and note that $\mathcal{S}_{\phi}^{1},\dots,\mathcal{S}_{\phi}^{k},\mathcal{S}_{\phi}\setminus\bigcup_{j=1}^{k}\mathcal{S}_{\phi}^{j}$ is a partition of $\mathcal{S}_{\phi}$. By Proposition 2.1 and Lemma 2.2(4),
$\sum_{j=1}^{k}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}=\sum_{j=1}^{k}\|\mathcal{S}_{\phi}^{j}\|_{s\text{-var}}\leq\|\mathcal{S}_{\phi}\|_{s\text{-var}}<\infty.$
Suppose now that $\sum_{j=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{j}))^{s}<\infty$. Let $X_{1},\dots,X_{n}$ be a partition of $\mathcal{S}_{\phi}$, where the arcs $X_{j}$ are enumerated according to the orientation of $\mathcal{S}_{\phi}$, with $0\in X_{n}$. We consider three subsets $P_{1},P_{2},P_{3}$ of the indices set $\{1,\dots,n\}$, and for indices in each $P_{i}$ we prove corresponding estimates.
Estimate 1: Set $P_{1}=\{n\}$, and let $k\in\mathbb{N}$ be the maximal index such that
$X_{n}\subset\overline{\mathcal{S}_{\phi}\setminus(\mathcal{S}_{\phi}^{1}\cup\dots\cup\mathcal{S}_{\phi}^{k-1})}.$
Then, by definition of $\mathcal{S}_{\phi}^{k}$ and decreasing property of their diameters, by Proposition 2.1 and by Lemma 2.2 we have
$(\operatorname{diam}{X_{n}})^{s}\leq(\operatorname{diam}\mathcal{S}_{\phi}^{k})^{s}\leq\|\mathcal{S}_{\phi}^{k}\|_{s\text{-var}}=(H_{1/s}(\mathcal{S}_{\phi}^{k}))^{s}.$
Estimate 2: Let
$L=\{l\in\{1,\dots,k\}:\,\text{there exists }j\in\{1,\dots,n\}\,\,\text{such that}\,\,X_{j}\subset\mathcal{S}_{\phi}^{l}\},$
and set
$P_{2}=\{j\in\{1,\dots,n\}:\,\text{there exists }l\in L\,\,\text{such that}\,\,X_{j}\subset\mathcal{S}_{\phi}^{l}\}.$
Given $l\in L$, let $\{j_{1}^{l},\dots,j_{m}^{l}\}$ be a maximal set of indices in $P_{2}$ such that $X_{j_{1}^{l}},\dots,X_{j_{m}^{l}}\subset\mathcal{S}_{\phi}^{l}$. Then, by Lemma 2.2, and by Proposition 2.1 we have
$(\operatorname{diam}{X_{j_{1}}})^{s}+\cdots+(\operatorname{diam}{X_{j_{m}}})^{s}$ $\leq\|X_{j_{1}}\|_{s\text{-var}}+\cdots+\|X_{j_{m}}\|_{s\text{-var}}$
$\leq\|X_{j_{1}}\cup\cdots\cup X_{j_{m}}\|_{s\text{-var}}$
$\leq\|\mathcal{S}_{\phi}^{l}\|_{s\text{-var}}$
$=(H_{1/s}(\mathcal{S}_{\phi}^{l}))^{s}.$
Estimate 3: Finally, set $P_{3}=\{1,\dots,n\}\setminus(P_{1}\cup P_{2})$. For each $j\in P_{3}$, there exists minimal $k_{j}\in\mathbb{N}$ such that $X_{j}\cap\mathcal{S}_{\phi}^{k_{j}}\neq\emptyset$. Moreover, if $j,j^{\prime}$ are two distinct such indices, then $k_{j}\neq k_{j^{\prime}}$. For any $j\in P_{3}$, by the decreasing property of $(\operatorname{diam}\mathcal{S}_{\phi}^{m})_{m\in\mathbb{N}}$, we have
$(\operatorname{diam}{X_{j}})^{s}\leq(\operatorname{diam}{\bigcup_{m=k_{j}}^{\infty}\mathcal{S}_{\phi}^{m}})^{s}\leq(\operatorname{diam}\mathcal{S}_{\phi}^{k_{j}})^{s}\leq\|\mathcal{S}_{\phi}^{k_{j}}\|_{s\text{-var}}=(H_{1/s}(\mathcal{S}_{\phi}^{k_{j}}))^{s}.$
Putting the three estimates together, we get
$\sum_{j=1}^{n}(\operatorname{diam}{X_{j}})^{s}=\sum_{i=1}^{3}\sum_{j\in P_{i}}(\operatorname{diam}{X_{j}})^{s}\leq 2\sum_{n=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{n}))^{s},$
since the term $(H_{1/s}(\mathcal{S}_{\phi}^{k}))^{s}$ may appear at most two times, through the sums over $j$ in $P_{1}$ and $P_{2}$, and the rest of the terms may appear at most twice through the sums over $P_{2}$ and $P_{3}$. Since the partition $\{X_{1},\dots,X_{n}\}$ is arbitrary, it follows that
$\|\mathcal{S}_{\phi}\|_{s\text{-var}}\leq 3\sum_{n=1}^{\infty}(H_{1/s}(\mathcal{S}_{\phi}^{n}))^{s}<\infty,$
and, by Proposition 2.1, $\mathcal{S}_{\phi}$ is a $(1/s)$-Hölder arc. ∎
4. Hölder exponents between spirals
Suppose $0<p\leq q$, and define the spiral arc
$\mathcal{S}_{p,q}=\{t^{-p}\cos t+\mathbf{i}t^{-q}\sin t:t\in[2\pi,\infty)\}\cup\{(0,0)\}.$
Burrell, Falconer, and Fraser in [BFF22, Theorems 2.9, 2.11] gave the following upper bounds on the Hölder exponent for maps between such spirals.
Theorem 4.1 ([BFF22]).
Suppose $f:\mathcal{S}_{p,q}\rightarrow\mathcal{S}_{r,s}$ is $\alpha$-Hölder, with $r\leq 1$. If $p>1$, then
(4.1) $\alpha\leq\frac{1+s}{2+s-r}.$
Otherwise, if $p\leq 1$, then
(4.2) $\alpha\leq\frac{p+q+r+s-pr+qs}{(2+s-r)(1+q)}.$
Theorem 1.1 also provides bounds on the Hölder exponent of maps between such spirals in an implicit way. Recall that there are functions $\phi,\psi:[2\pi,+\infty)\rightarrow(0,\infty)$ that tend to 00 as $t\rightarrow\infty$ with $\mathcal{S}_{p,q}=\mathcal{S}_{\phi}$ and $\mathcal{S}_{r,s}=\mathcal{S}_{\psi}$. Note that it is non-trivial to explicitly determine $\phi$ and $\psi$, due to the implicit relation between arguments of
$z_{t}=t^{-p}\cos t+\mathbf{i}t^{-q}\sin t\in\mathcal{S}_{\phi},$
for some $t\geq 2\pi$, and the modulus $|z_{t}|$. Namely, while the distance of $z_{t}$ from 00 is indeed just $|z_{t}|$, the value $t$ is not always an argument of $z_{t}$, which makes the naive approach of choosing $\phi(t)=|z_{t}|$ generally incorrect for $p\neq q$. However, it is easier to determine $\phi(t_{k})$ and $\psi(t_{k})$ at $t_{k}=k\pi/2$, for integers $k\geq 4$, which is enough to imply that $\mathcal{S}_{\phi},\mathcal{S}_{\psi}$ are almost circular. This can also be seen through the relation of these spirals to the corresponding concentric ellipses (see [BFF22, p. 7]). Moreover, by the aforementioned values $\phi(t_{k})$ and $\psi(t_{k})$, we conclude that
$\phi_{j}=(2\pi j)^{-p},\,\,\,\psi_{j}=(2\pi j)^{-r},$
for all $j\in\mathbb{N}$. Suppose that $h:\mathcal{S}_{p,q}\rightarrow\mathcal{S}_{r,s}$ is $\alpha$-Hölder. By Theorem 1.1, there is a $\beta$-Hölder map $g:[0,1]\rightarrow\mathcal{S}_{p,q}$, for all $\beta<p$. Thus, $h\circ g:[0,1]\rightarrow\mathcal{S}_{r,s}$ is $\alpha\beta$-Hölder, which by Theorem 1.1, and the fact that $\beta$ can be as close to $p$ as needed, implies that
(4.3) $\alpha\leq\frac{r}{p}.$
Suppose $p>1$ and $r\leq 1$. The latter implies that
$\frac{(2-r)r-1}{1-r}<r\leq s,$
which leads to
$\frac{r(2-r+s)}{1+s}\leq 1\leq p.$
The above inequality is enough to conclude that
$\frac{r}{p}\leq\frac{1+s}{2+s-r},$
which shows that the inequality (4.3) achieved by Theorem 1.1 is indeed an improvement upon the bound (4.1).
The bound (4.3) is also an improvement on the bound stemming from Theorem 4.1 for polynomial spirals with $p=q\in(1,+\infty)$ and $0<r=s\leq 1$, i.e., an improvement on [BFF22, Corollary 2.10]. In fact, this is a sharp bound even for more general positive $p=q$, $r=s$, as stated in Theorem 1.2, which we are ready to prove.
Proof of Theorem 1.2.
Let $0<r\leq p$. If $h:\mathcal{S}_{p}\rightarrow\mathcal{S}_{r}$ is $\alpha$-Hölder, then is has already been shown that $\alpha\leq p/r$ in (4.3).
The desired maps between the spirals $\mathcal{S}_{p},\mathcal{S}_{r}$ are in fact appropriate radial stretch maps. In particular, the map $f:\mathbb{C}\rightarrow\mathbb{C}$ defined by
$f(z)=|z|^{\frac{r}{p}-1}z$
for all $z\neq 0$, and $f(0)=0$, is $r/p$-Hölder (see, for instance, [V7̈1, p. 49] and [AIM09, Corollary 3.10.3]) and satisfies $f(\mathcal{S}_{p})=\mathcal{S}_{r}$. Moreover, define the map $g:\mathbb{C}\rightarrow\mathbb{C}$ by $g(0)=0$ and
$g(z)=|z|^{\frac{p}{r}-1}z,$
for all $z\neq 0$. This map satisfies $g(\mathcal{S}_{r})=\mathcal{S}_{p}$ and is Lipschitz, due to the derivative being bounded on the closed disk $D(0,(2\pi)^{-r})$. This completes the proof. ∎
It should be noted that the improved bound (4.3) in the context of Hölder classification for spirals $\mathcal{S}_{p}$, $\mathcal{S}_{r}$ coincides with the sharp bound in the quasiconformal classification problem resolved by Tyson and the first author in [CGT23]. In particular, by [CGT23, Theorem 1.1], two spirals $\mathcal{S}_{p}$ and $\mathcal{S}_{r}$ are $K$-quasiconformally equivalent if, and only if, $K\geq p/r$ (see [V7̈1] for more details on quasiconformal mappings). It is quite interesting that in the case of these spirals, the sharp exponent bound in the Hölder classification programme suggested by Fraser in [Fra21] is essentially attained from the sharp dilatation bounds for the corresponding quasiconformal classification study. This motivates further the question of whether and under what conditions resolving the quasiconformal classification problem for two objects results in the resolution of the corresponding Hölder classification problem. We refer the interested reader to the discussion in [CG25, Section 5] for more details and related results in higher dimensions.